T(-10,-2) R(10,8) Y(11,-9) are vertices of ∆TRY
1) find equation of YA, the altitude from Y to TR . show that A(4,5)
2) if equation of perpendicular bisector is y=3x-7, find equation of circle through T, R, Y .
3) find the area of ∆TRY
4) find size of angle T

Answer by mananth(16949)

Equation of TR
x1 y1 x2 y2
-10 -2 10 8

slope m = (y2-y1)/(x2-x1)
( 8 - -2 )/( 10 - -10
( 10 / 20 )
m= 0.50

Plug value of the slope and point ( -10 , -2 ) in
Y = m x + b
-2.00 = -5 + b
b= -2.00 - -5
b= 3
So the equation of TR will be
Y = 1/2 x + 3 ( m=1/2)..................................(1)
AY is perpendicular to TR y=(11,-9)
slope of AY = -2 ( negative reciprocal)
Equation AY
slope =2 , Y=(11,-9)
(y-(-9))= (-2)(x-11)
y+9= -2x+22
y= -2x+13.................................................(2)
Solve (1)&(2) to get point A (4,5)
.
--------------------